(show equations used)

A 12.0 V battery is hooked up to two resistors in series: a 1000 ohm resistor and a 2200 ohm resistor.

A) Calculate the voltage across each resistor.

B) Calculate the current through each resistor.

C) What is the total resistance of the circuit?
Answer #1

(c)

Total resistance R = R1+R2 = 1000+2200 = 3200 ohm

(B) I = V/R = 12/ 3200 .= 3.75mA

(A) voltage across each resistor is ,

V1 = I*R1 = 0.00375* 1000 = 3.75V

V2= I* R2 = 0.00375 *2200 = 8.25V

I hope it was helpful.. please rate me positive..